# Synthetic pulsar lightcurves from global kinetic simulations and comparison with the Fermi-LAT catalog

### v1

@@ 3.2 Emission pattern and geometric origin of pulses
where $F$ is the photon flux received by a given observer. This quantity corrects for the anisotropy of the emission to infer the true luminosity. It is of the order $f_{\Omega}\sim 1$ in the equatorial regions where most of the emission comes from, and it quickly rises for any observer looking at the polar regions where little emission is predicted. The emission pattern presents a bright sinusoidal-like component (or “caustic”) at low obliquities which breaks up into two bright hotspots at higher obliquities ($\chi\gtrsim 30^{\rm o}$). The skymaps present a central symmetry with respect to the center of the maps, located here in the equatorial plane at the pulsar phase $\Phi=0.5$. Slight asymmetries between the left and the right caustics can be seen in some solutions like in the $\chi=45^{\rm o}$ solution; these are due to statistical fluctuations and we anticipate that they would vanish for a longer integration time. A pulse of light is received by a distant observer each time the current layer passes through the line of sight, which happens in general twice per pulsar period. The central symmetry in the emission maps is due to the periodic change in the direction of the current density direction carried by the layer (mostly oriented towards ±$\pm\theta$ directions) between two consecutive crossings, leading to a preferential emission in the northern/southern hemispheres in each hotspot.
@@ 3.5 Pulse width
This work constrains the location of the pulsed TeV emission discovered in the Crab and Vela pulsars within the inner parts of the pulsar wind. The target photon field must therefore remain compact. A promising candidate for emitting this low-energy radiation are secondary pairs produced in the vicinity of the current sheet –like a sheath surrounding the layer– near the light cylinder and shinning synchrotron radiation (Lyubarskii, 1996; Hakobyan et al., 2019). This photon field would appear isotropic seen from the high-energy pairs accelerated deep inside the reconnection layer, which could then be upscattered into the TeV range. In contrast, any diffuse target photon field from the surrounding nebula, the CMB, or from a nearby main-sequence companion star would smear out any pulsation.
@@ 5 Summary
  • The current layer is not uniformly bright at all latitudes, it is mostly shinning in the equatorial plane. As a direct consequence, pulsars with nearly symmetric pulses separated by $0.5$ in phase should present the highest apparent luminosities.
@@ Acknowledgements.
We thank Alexander Philippov, Alexander Chernoglazov, Christo Venter, and Constantinos Kalapotharakos for useful comments regarding this work. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 863412). Computing resources were provided by TGCC under the allocation A0130407669 made by GENCI and PRACE (call 17).

### v2

@@ 3.2 Emission pattern and geometric origin of pulses
where $F$ is the photon flux received by a given observer. This quantity corrects for the anisotropy of the emission to infer the true luminosity. It is of the order $f_{\Omega}\sim 1$ in the equatorial regions where most of the emission comes from, and it quickly rises for any observer looking at the polar regions where little emission is predicted (see also Kalapotharakos et al. 2023). The emission pattern presents a bright sinusoidal-like component (or “caustic”) at low obliquities which breaks up into two bright hotspots at higher obliquities ($\chi\gtrsim 30^{\rm o}$). The skymaps present a central symmetry with respect to the center of the maps, located here in the equatorial plane at the pulsar phase $\Phi=0.5$. Slight asymmetries between the left and the right caustics can be seen in some solutions like in the $\chi=45^{\rm o}$ solution; these are due to statistical fluctuations and we anticipate that they would vanish for a longer integration time. A pulse of light is received by a distant observer each time the current layer passes through the line of sight, which happens in general twice per pulsar period. The central symmetry in the emission maps is due to the periodic change in the direction of the current density direction carried by the layer (mostly oriented towards ±$\pm\theta$ directions) between two consecutive crossings, leading to a preferential emission in the northern/southern hemispheres in each hotspot.
@@ 3.5 Pulse width
3.6 Implications for the pulsed TeV emission
This work constrains the location of the pulsed TeV emission discovered in the Crab and Vela pulsars within the inner parts of the pulsar wind. The target photon field must therefore remain compact. A promising candidate for emitting this low-energy radiation are secondary pairs produced in the vicinity of the current sheet –like a sheath surrounding the layer– near the light cylinder and shining synchrotron radiation (Lyubarskii, 1996; Hakobyan et al., 2019). This photon field would appear isotropic seen from the high-energy pairs accelerated deep inside the reconnection layer, which could then be upscattered into the TeV range. In contrast, any diffuse target photon field on much larger scales, from the surrounding nebula, the CMB, or from a nearby main-sequence companion star, would smear out any pulsation by adding up a dominant contribution from large radii.
The detection of 20 TeV pulsed emission from the Vela pulsar implies that pairs must be accelerated up to at least $\gamma_{\rm max}\gtrsim 4\times 10^{7}$ (H. E. S. S. Collaboration et al., 2023). This is a challenging constraint for the models because this energy scale is significantly above the radiation-reaction-limited particle energy estimated at the light cylinder, $\gamma^{\rm LC}_{\rm rad}\approx 4\times 10^{5}\ll\gamma_{\rm max}$ (assuming $E=\tilde{B}_{\perp}=B_{\rm LC}\approx 5\times 10^{4}$, see Eq. 7). One possible explanation for this discrepancy involves X-point acceleration, where radiative losses are substantially reduced, meaning that particle acceleration could continue nearly unimpeded up to the upstream magnetization limit, $\gamma_{\rm max}\gtrsim\sigma_{\rm LC}$ (Cerutti et al., 2012a; Kagan et al., 2016; Chernoglazov et al., 2023). This would in turn imply that the magnetization must be much higher than usually assumed in Vela.
Another effect that could alleviate this issue is particle reacceleration at larger radii, as reported in Sect. 3.5, because reconnection pursues magnetic dissipation while radiative losses diminish thanks to the weakening of the magnetic field strength with distance. Note that the spatial offset between the peak of the synchrotron and the inverse Compton emission would not lead to any significant changes in pulse phase, thanks to the caustic effect highlighted in Sect. 3.2. However, it remains uncertain whether this effect alone can account for the required hundred times increase in particle energy, from $\gamma_{\rm rad}\approx 4\times 10^{5}$ to $\approx 4\times 10^{7}$, on reasonably compact scales (i.e., $\hat{r}\lesssim r_{\rm m}$). Despite these challenges, a combination of both mechanisms may render this scenario viable, but further investigation is necessary to explore these questions in greater detail.
@@ 5 Summary
  • The current layer is not uniformly bright at all latitudes, it is mostly shining in the equatorial plane. As a direct consequence, pulsars with nearly symmetric pulses separated by $0.5$ in phase should present the highest apparent luminosities.
@@ Acknowledgements.
We thank Alexander Philippov, Alexander Chernoglazov, Christo Venter, and the referee Constantinos Kalapotharakos for useful comments regarding this work. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 863412). Computing resources were provided by TGCC under the allocation A0130407669 made by GENCI and PRACE (call 17).